Sherman is 4 times as old as his pet rattlesnake. In 5 years, assuming they are both still around, Sherman's age will be 3 years more than twice the snake's age. How old is each?
Answer by JulietG(1812)

S = 4R [Sherman is 4 times as old as his pet rattlesnake.]
S + 5 = 2(R+3) + 3 [In 5 years, assuming they are both still around, Sherman's age will be 3 years more than twice the snake's age.]
Substitute the value of S in the first equation into the second.
(4R) + 5 = 2(R+3) + 3
Distribute
4R + 5 = 2R + 6 + 3
Add
4R + 5 = 2R + 9
Subtract 2R from each side
2R + 5 = 9
Subtract 5 from each side
2R = 4
Divide each side by 2
R = 2
.
If the rattler is 2, then Sherman is 4*2, or 8.
.
Let's plug the values into the second equation to make certain.
S + 5 = 2(R+3) + 3
8 + 5 = 2(2+3) + 3
13 = 2(5) + 3
13 = 10 + 3
13 = 13
Success!
.
No one should keep a rattler as a pet. They're wild animals.
They sure are delicious when cooked over a spit, though! "Mmm, tastes like chicken!"